40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog, Page 42 of 58  Not logged in ELOG logo
ID Date Author Type Category Subject
  821   Tue Aug 13 14:24:11 2013 EvanLaserCavityCavity skewness and spot locations on windows

 I've taken the above displacements and angles and used them to project the cavity spots onto the cryostat windows, assuming the radiation shields can be perfectly aligned to the window axes.

In both the Solidworks model and the actual cryostat, the radiation shields are displaced inwards relative to the window axes. Dmass has measured the distance of closest approach of the rad shields to be 1/2 inch. In the Solidworks model, this means that the west rad shield is displaced 0.31 inches toward the east relative to the west windows, and the east rad shield is displaced 0.31 inches toward the west relative to the east windows. I've included these displacements in the following calculations.

I used two different methods to project the spot locations: (1) Solidworks, and (2) python code. For Solidworks (see first attachment for picture), I first aligned the radiation shields to the window axes (plus the aforementioned displacements), and then used the move/rotate tool to skew the cavities. I then projected the cavity axes onto the windows and used the measure tool to get the displacement between the spot and the window center. The python code (attached) is fairly simple; it's just takes the displacements and angles and propagates them using the small angle approximation. Since the windows are at 19° relative to the cavity faces, there's an obliquity factor of cos(19°) that I correct for in the horizontal coordinates.

The spot locations as determined from the Solidworks drawing are as follows:

  West cavity East cavity
frontX +256 mil −468 mil
frontY −75 mil −124 mil
backX −387 mil +217 mil
backY −85 mil −41 mil

The spot locations as determined from my python code are as follows:

  West cavity East cavity
frontX +253±7 mil −462±8 mil
frontY −79±6 mil −125±8 mil
backX −374±7 mil +222±8 mil
backY −87±6 mil −43±8 mil

The conclusion that Dmass and I have drawn from this is that it is not necessary to engineer any pitch correction into the supports for the radiation shields.

Attachment 1: dewar_and_cavity_stripped_skewed.jpg
dewar_and_cavity_stripped_skewed.jpg
Attachment 2: windowSpots_code.zip
  820   Mon Aug 12 16:48:32 2013 Jonathan MishlerComputing Frequency Comb "Heat Plot"

This is only for one specified comb repetition frequency.

[dym: way too many inputs and outputs here - make it have one output, and two inputs. If you want more info/control later, you can add stuff, but keep it simple while you go through and debug.]

Attachment 1: FreBeat.m
function [LongCombFrequencies, ShortCombFrequencies, CombBeats, MinBeat, MinBeatRow, MinBeatColumn, CombTeethNumberSum] = FreBeat(LongCombMaxFrequency, AdjustedLongCombCEO, LongCombREP, ShortCombMaxFrequency, AdjustedShortCombCEO, ShortCombREP)

% The function FreBeat combines two frequency combs and calculates every
% single beat between every comb tooth. Once every beat is calculates, the
% smallest beats, and the combs that give the smallest beats are
% identified.

% The maximum frequencies, repetition frequencies, and adjusted carrier
% envelope offset (CEO) frequencies are the function's input. For
% clarification, the adjusted CEO frequencies are simply each comb's
... 46 more lines ...
  819   Mon Aug 12 15:34:43 2013 DmassCryostatCavityNew Cavity Suspension - springs

Choosing a spring for the platform cavity suspension: 

 ===========================

      BeCu properties:    

===========================

I hunted some for the NIST cryogenic BeCu property reference (the existence of which Warren hinted at), but all I found was this NIST page on other cryogenic BeCu properties, there were no mechanical properties here

  • I found a Springer book: "Low Temperature Material Properties" (Table 2.8)
    • Table 2.8 gives the following values for E_y for Cu + 2Be:
      • E_y(0 K) = 134 GPa
      • E_y(80 K) = 130 GPa
      • E_y(300 K) = 118 GPa
  • I found a publications by the Copper Development association titled "Beryllium Copper" which sounded like it might have relevant info...
    • Table 5 gives values for the Young's modulus of BeCu for the following prep conditions (yes they report in gigapounds per square inch):
    • Solution heat treated and precipitation hardened:
      • E(300K) = 17.6 Glb/in^2
      • E(250K) = 17.9 Glb/in^2
      • E(200K) = 18.2 Glb/in^2
      • E(150K) = 18.8 Glb/in^2
      • E(100K) = 19.2 Glb/in^2
    • Solution heat treated, cold drawn to "half hard" condition and precipitation hardened:
      • E(300K) = 18.11 Glb/in^2
      • E(250K) = 18.5 Glb/in^2
      • E(200K) = 18.5 Glb/in^2
      • E(150K) = 18.4 Glb/in^2
      • E(100K) = 19.2 Glb/in^2
    • I don't know what the prep conditions of "normal" BeCu are, and if either of these are representative of that. Warren might be able to help 
  •  In this giant beast of a document from NIST (scanned by some folks @ the Princeton Plasma Physics Lab):
    • Table 13.1 (~ page 629) shows a huge variation in BeCu Young's moduli based on prep method (115 to 172 GPa at room temperature)
    • Table 13.2 shows three some cryogenic data for "Cold-worked and aged C17200 BeCu":
    • E(227K) = 124 GPa
    • E(261K) = 122 GPa
    • E(297K) = 120 GPa
  • Warren's memory said that the increase in Young's modulus was ~10% when going from 300K to 120K. I will use this number until I hear otherwise / find something I am more sure of than that.

 =================================

    How do springs work?

==================================

It's worth writing down the super basics of how springs work so that I can connect these real world numbers to the idealized freshman physics numbers I want to use on an envelope in the design/choice of springs.

To first order, I will assume:

  •  Lo(300K) ~ Lo(120K) 
  • 1.1*k(300K) ~  k(130K)
  • Where Lo is the unstretched length of the spring, and k is it's spring constant.

I found this spring constant calculator, which lets you plug in length, diameter, wire material, and wire thickness, and gives you a bunch of paremeters about the spring, and lets you submit whatever parameters you enter to request a quote for said spring.

They use the following equation for the extension spring constant:

k = G d^4 / (8 D^3 N)

  • d = wire size 
  • D = mean diameter (diameter - not outer diameter - of spring. D = OD - d)
  • N = # of coils
  • G = Shear modulus of material

D/d = 10 seems like a very reasonable place to start (probably within a factor of 2 of the right ratio)

 =================================

    What spring do I want?

==================================

We need to be slightly careful about how we choose the spring and design the system because of the changing Young's modulus of BeCu...

  • Spring has unstressed length Lo
  • Spring has stressed length Lo + mg/k, where m is (1/4) the mass of the platform + cavity + shield assembly
  • We have some freedom in choosing Lo and k; there is a continuum of solutions
  • The longer we make the spring, the stiffer it has to be, and the higher the resonant frequency of vertical isolation (seems bad?)
  • The Longer we make the spring, the smaller mg/k is, and thus the LESS the height of the cavity will change when we cool down the assembly.

The windows are ~31/32" clear aperture vertically, though we want to stay within the inner 1/2" if possible

How low can I put the platform?

It seems like we should be able to have the platform *not* touch the bottom of the inner can at 300K, especially because it rattling around in there while we raise and lower it will probably kick the cavities around and prevent us from ever guaranteeing that they stay in the same place that we place them in. Having the top of the platform at 5.875" puts the beams through the middle of the window, and gives us ~3/16" clearance between the bottom of the (Thorlabs 1/2") platform and the bottom of the inner can at room temperature. We will use this height.

How does Lo relate to the vertical raising of the platform when we cool the suspension?

  • Top of platform is 5.875" below the cold plate
  • The obvious limiting case: Lo = 0 implies
    • mg/k = 5.875" when warm
    • mg/k => mg/(1.1k) = 5.34" when cold, which would put the beam way too close to the top of the window
    • height increase = 5.875 - 5.34 = 0.535"
  • Lo = 2.9" => height increase = ...
    • (5.875 - 2.9) - (5.875 - 2.9)/1.1 = 0.27"
  • Etc...

This doesn't seem impossible yet!

 

  818   Fri Aug 9 15:50:27 2013 EvanLaserCavityCavity skewness relative to outer rad shields

Last week Dmass and I measured the positions of the cavity axes relative to the outer radiation shield axes. We aligned the beams to the cavity axes. At each aperture, Dmass held a card flush with the face of the shield to make the spot visible, and I took some photographs.

For each of the back (output) apertures, the camera was positioned looking dead-on into the aperture. For each of the front (input) apertures, the camera was placed in two different positions: vertically above the beam axis, and horizontally to the side of the beam axis. For each camera position I took two photos.

I opened each photo in Inkscape and drew an ellipse which (by eye) coincided with the edge of the aperture (representative photo attached; the full set is in the ligo.wbridge picasa album). The major and minor axes of the ellipse are constrained to lie horizontally and vertically, so here we're exploiting the assumption that (a) the roll of the camera was negligible, and (b) the oblique viewing angle of the camera was either entirely horizontal or entirely vertical for each photograph. I read off the coordinates of its four vertices (top, bottom, left, and right) in terms of pixels, as well as the center of the beam spot (determined by eye). Since the aperture is a circle with 1/2-inch diameter, I used the horizontal axis of the ellipse (in pixels) to convert the horizontal coordinate of the beam spot in pixels to the horizontal coordinate in inches from the center of the aperture, and likewise for the vertical coordinate. I also assigned uncertainties by eye and propagated them forward.

To maintain a consistent, right-handed 3D coordinate system, the horizontal coordinates for the back face measurements are given a sign flip. Then +z is normal to and directed outward from the front face, +x points from west to east, and +y points upward. The shield apertures are separated by z = 6.5 inches. To get the displacement of the cavity axis relative to the shield axis, we take the average (xFront + xBack, yFront + yBack) / 2. To get the pitch and yaw of the cavity axis, we take the (x, y, z) coordinates of the two spots, subtract the back coordinates from the front coordinates, normalize the resulting vector, and then read off (pitch, yaw, 1). The resulting angle convention is that positive pitch means the back of the cavity is tilted up relative to the front of the cavity (positive rotation about the x axis), and positive yaw means that the cavity is rotated counterclockwise when looking downward (positive rotation about the y axis).

  West cavity East cavity
x displacement −14±2 mil −13±3 mil
y displacement −83±2 mil −84±3 mil
yaw −0.39±0.04 deg −0.77±0.05 deg 
pitch -0.03±0.04 deg  +0.28±0.05 deg

 

[dym (adding data for completeness, and complained about convention for pitch being backwards so that we fixed it).

In order to be careful while we were doing these measurements, I used a small metal ruler to measure the beam spot location with respect to the outer radiation shield aperture, and generate the same numbers that Evan did using the pictures we took. All errors were set to be 1/128", which is 1/4 of the 1/32 scale I was using to do the measurements (a.k.a. "this is how well I think I could identify the spot location w.r.t. the edges of the radiation shield using the ruler and my eye")

West Cavity East Cavity
x displacement -12 ± 11 mil -31 ± 11 mil
y displacement -78 ± 11 mil -86 ± 11 mil
yaw -0.41 ± 0.1 deg -0.83 ± 0.1 deg 
pitch 0.0 ± 0.1 deg 0.41 ± 0.1 deg

Once we use conventions for pitch that are consistent with the coordinate system we agreed upon, then our numbers agree in sign, and are close in value w.r.t. the error bars. Only the East Cavity pitch seems to disagree, but since the numbers for the pictures seemed better all around (by error estimates), we are using these for the shield redesign]

Attachment 1: crop_east_back_1_annot.jpg
crop_east_back_1_annot.jpg
Attachment 2: skew_code.zip
  817   Thu Aug 8 17:52:30 2013 DmassLaserLab WorkCavity scans + mode matching

 We didn't have a recorded quantitative measure of the mode mismatch of the cavities (though we had claimed that it was as bad as 40% based on transmitted power in the East cavity.

I plugged the transmitted and reflected PD signals into the scope (these have been realigned since we opened up the cryostat and realigned/relocked in air).

I drove the current driver (1mA/V) with a function generator (3Hz, 2 Vpp) T-ed into the scope (Zin = 1M).

Turned down the total power at the refl PD by adjusting the HWP at the laser output before the isolator to stop saturating the DC path of the PD (P_incident = 2mW)

I tweaked the alignment to maximize the transmission / reflection dips

I turned down the speed / voltage range while looking at the dips in reflection on resonance the peaks (transmission and reflection) stopped getting bigger (because at high enough speeds, I don't fill the cavity).

Sweeps:

East:

Are we sweeping slow enough?:

  • Sweep speed: [ 2 V / (1 / 6 s) ]  x  [ 1 mA / V ]  x  [ 23.7 MHz / mA ]   =   280 MHz / sec
  • Cavity pole ~ 30kHz
  • t_fill ~ 1 / f_pole = 33 us
  • t_sweep = 30kHz / [ 280 MHz / sec ] = 107 us
  • t_sweep ~ 3 x t_fill (so we can treat the sweeps like steady state measurements)

Reflection:

  • Vmax = 223 mV
  • Vmin = 45 mV
  • Vdark = - 20.6 mV (* yes this is large, idk why)
  • % not reflected (trans + loss) = ( ( Vmax - Vdark ) - ( Vmin- Vdark ) ) / ( Vmax - Vdark ) * 100 = (Vmax - Vmin) / (Vmax - Vdark)*100 = (223 - 45) / (223 + 20.6)*100
  • West trans + loss = (223 - 45) / (223 + 20.6)*100 = 73%***

West:

** I noticed that while sweeping with the same magnitude sweep that I used for aligning the East cavity (10Vpp and 10Hz), the dips in reflection/transmission seem to jump around a lot more than they did with the East cavity. I turned off the laser current and temperature, waited for 5 minutes, and then turned back on temp, waited 5 minutes, and turned on the current. It's possible that one of the temperature loops is less stable than the other (since they are little plastic screw knobs for PID tuning with no readback on two different ITC510 controllers. The oscillations were ~1s timescale and the thermal pole for the TEC/laser is ~6Hz, so this might not be crazy.

 Are we sweeping slow enough?:

  • Sweep speed: [ 2 V / (1 / 6 s) ]  x  [ 1 mA / V ]  x  [ 38.8 MHz / mA ]   =   466 MHz / sec
  • Cavity pole ~ 30kHz
  • t_fill ~ 1 / f_pole = 33 us
  • t_sweep = 30kHz / [ 466 MHz / sec ] = 65 us
  • t_sweep ~ 2 x t_fill (so we can treat the sweeps like steady state measurements)

Reflection:

  • Vmax = 200 mV
  • Vmin = 32.4 mV
  • Vdark = - 1.6 mV
  • % not reflected (trans + loss) = (Vmax - Vmin) / (Vmax - Vdark)*100 = (200 - 32.4) / (200 + 1.6)*100
  • East trans + loss = (200 - 32.4)/(200 + 1.6) = 83%***
  816   Thu Aug 8 14:09:33 2013 DmassDailyProgress Overcoming a bandwidth limited frequency comb

Quote:

Although designing a microresonator with ~10 GHz mode spacing is not difficult, there is the problem that as the diameter increases, the four wave mixing process is largely suppressed (maybe some group has fixed this). For octave spanning frequency combs generated with microresonators, a tunable ECDL (1550 - 1620 nm) is used. As seen in literature, an 80 micro meter diameter silica microresonator will produce a comb with 850 GHz spacing. If the central wavelength used is 1550 nm, there will be a comb tooth < 10 GHz away from 1064 nm. So the 1064 nm laser can easily be locked to this comb. Once the 1064 nm laser is locked to this comb, the stability of the 1064 nm laser will be transferred to each comb tooth. An f-2f interferometer can then be used to fully stabilize the 850 GHz comb. 

My idea: If each comb tooth now has the stability of the 1064 nm laser, why can't we lock one of those comb teeth to a tooth from a new comb with 10 GHz spacing that has a bandwidth from 1540 - 1560 (this comb exists, I have seen it in literature). By doing this, we will have transferred the stability of the 1064 nm source to each comb tooth (10 GHz comb). We wouldn't need an f-2f interferometer to stabilize the the 10 GHz comb. If we combine two frequency combs, one of which is fully stabilized, and one of which we know the repetition frequency, it is simple math to calculate to CEO frequency of the other comb. Once a computer calculates the CEO frequency of the other comb, it can completely stabilize the 10 GHz comb. 

Once the 10 GHz comb is stabilized, all we have to do is beat a comb tooth with the 1550 nm laser.

I am wondering if this seems reasonable?

 

The tl;dr is: "this won't work as stated, BUT we might be able to do something slightly more clever than this"

One of the assumptions implicit in what you stated is that 1064 and 1550 are not the EXACT wavelengths of the lasers down in lab that we're trying to lock to. This is untrue. To be more quantitative do the following:

  • Calculate how many nm 10GHz corresponds to:
    • 10GHz * (1550 nm/1.9e14 Hz) = 8.2e-11 m
    • 10GHz * (1064 nm/2.8e14 Hz) = 3.8e-11 m
  • Ask "how many digits of precision does this imply we need to know the 1064nm and 1550nm wavelengths to?"
    • 8.2e-11 m/1550nm = 5.3e-5
    • 3.8e-11 m/1064nm = 3.6e-5
  • Are the lasers downstairs actually this close to those wavelengths?
    • 1550.00 +/- 0.08nm
    • 1064.00 +/- 0.04nm

Now that we have arrived at the correct question, it looks grim. The 1550 nm diodes are in ITU C-band 34, which means they have a ~1550.12 +/- 0.4nm. I do not know what exactly comes out of the NPROs as built (the 1064nm lasers in Tara's lab).

HOWEVER:

IF we only have to stabilize the repetition rate (which controls the spacing) instead of needing it at some specific value, then we ***might*** be able to make a tunable comb.

The questions we would need to answer:

What is the range of tuning for the whatever we use to actuate on the carrier envelope offset (how many Hz can this move in some physical system?)

What is the range of tuning that we can get for the rep rate, how many lines are there, and what frequency range does THIS correspond to? (The hope is that we get an N-fold increase in the tuning range here, so if we tune the rep rate by 10GHz, and there are N lines between the ~1064nm line and the ~1550nm line, then the line by the 1550 nm line moves by N*10GHz.

We also may have a SLOW knob on the comb spacing in the temperature of the microresonator, which (along with the CTE for whatever the resonator is made out of), might give us a large range slow actuator. This would need to be calculated.

This may lead to just a diagonalization controls problem., which the nice LIGO people have enough experience with to not be so scary.

[DYM edit: IMPORTANT TAKE AWAY POINT:

It *seems like* your idea requires that the 1064nm laser in Tara's lab (we can measure this wavelength but don't assume it's 1064.000nm) be near a comb line in a 850 GHz spacing resonator. We briefly touched on this at last weeks meeting with Dave Reitze, and decided that it sounded like this requires you to fabricate the resonator to mechanical tolerances that allow you to put a line next to 1064nm, which sounds like it is really hard to do without making a bunch of microresonators and cherry picking the one you want. "How hard is it to build a microresonator with exactly 850GHz spacing and a center frequency of exactly 1550nm? How sensitive are the tolerances? Maybe we can tune the diameter of the microresonator enough by changing its temperature to overcome to imprecision inherent in the machining process?]

  815   Wed Aug 7 11:06:48 2013 Jonathan MishlerDailyProgress Overcoming a bandwidth limited frequency comb

Although designing a microresonator with ~10 GHz mode spacing is not difficult, there is the problem that as the diameter increases, the four wave mixing process is largely suppressed (maybe some group has fixed this). For octave spanning frequency combs generated with microresonators, a tunable ECDL (1550 - 1620 nm) is used. As seen in literature, an 80 micro meter diameter silica microresonator will produce a comb with 850 GHz spacing. If the central wavelength used is 1550 nm, there will be a comb tooth < 10 GHz away from 1064 nm. So the 1064 nm laser can easily be locked to this comb. Once the 1064 nm laser is locked to this comb, the stability of the 1064 nm laser will be transferred to each comb tooth. An f-2f interferometer can then be used to fully stabilize the 850 GHz comb. 

My idea: If each comb tooth now has the stability of the 1064 nm laser, why can't we lock one of those comb teeth to a tooth from a new comb with 10 GHz spacing that has a bandwidth from 1540 - 1560 (this comb exists, I have seen it in literature). By doing this, we will have transferred the stability of the 1064 nm source to each comb tooth (10 GHz comb). We wouldn't need an f-2f interferometer to stabilize the the 10 GHz comb. If we combine two frequency combs, one of which is fully stabilized, and one of which we know the repetition frequency, it is simple math to calculate to CEO frequency of the other comb. Once a computer calculates the CEO frequency of the other comb, it can completely stabilize the 10 GHz comb. 

Once the 10 GHz comb is stabilized, all we have to do is beat a comb tooth with the 1550 nm laser.

I am wondering if this seems reasonable?

 

Attachment 1: Screen_Shot_2013-08-06_at_3.17.47_PM.png
Screen_Shot_2013-08-06_at_3.17.47_PM.png
Attachment 2: MicroresonatorSize.xlsx
  814   Thu Aug 1 15:45:51 2013 Jonathan MishlerDailyProgressSchematicsUpdated Schematic and Notes

 In the schematic:

An external cavity diode laser is fed through an erbium doped fiber amplifier, filtered, sent through a polarization controller and into a microresonator by means of evanescent coupling through a fiber taper. The beam is then split. The comb is phase locked to a 1064 nm optical source, transferring its coherence to each comb tooth. The error signal drives the fiber tapering control system.

A portion of the comb is also sent to an f-2f interferometer. The output signal is fed and referenced to to an atomic oscillator. The repetition rate of the ECDL produces a signal which is fed to the atomic standard. These two signals being fed into the atomic standard stabilize the output of the ECDL.

The last comb output is beated with the 1550 nm source which compares the stability of the 1550 and 1064 sources.

Attachment 1: Screen_Shot_2013-07-31_at_2.15.04_PM.png
Screen_Shot_2013-07-31_at_2.15.04_PM.png
Attachment 2: Screen_Shot_2013-08-01_at_3.39.46_PM.png
Screen_Shot_2013-08-01_at_3.39.46_PM.png
Attachment 3: 13_notes.docx
Attachment 4: Microresonator_Literature.zip
  813   Thu Jul 25 16:11:12 2013 DmassCryostatCavityNew Cavity Suspension

Show drawings of suspension to Warren - mechanical or pictures - to refresh his memory about what this looks like

Warren: good data on change in young's modulus in BeCu @ Cryo temperatures - there is a website connected with the cryo division of NIST - dmass will hunt for this and check back with Warren if he has trouble

Search for: Cryo properties data site

Rana suggests directly mounting the cavity and the radiation shields to the platform individually.

Dmass is working to get measurements of whether or not the cavity axis is skew w.r.t. the Si spacer

Intensity noise coupling - Evan + Dmass set requirements for this

Eventually want to directly measure the intensity coupling, but intensity noise couples through the PDH loops in a nonlinear fashion - this will be tricky

NDS2 now works - set this up in the lab - Evan will do this

Dmass + Rich are going to get the transistor based EOM amplifier in play int he Cryolab

  812   Wed Jul 24 13:55:34 2013 Jonathan MishlerDailyProgressSchematicsFirst go at frequency comparison schematic

[Dmass, Jonathan]

Here is a very rough first shot at a schematic of a system that directly compares the stability of 1550 and 1064 nm sources.

 

Screen_Shot_2013-07-24_at_1.31.33_PM.png

In this figure, a tunable cw laser is amplified with Yb fiber, filtered, sent through a polarization controller, and fed through a silica microtoroidal resonator. A 1064 nm source is phase locked to the frequency comb generated by the microresonator. A computer tunes the cw laser and fiber taper to restrict the modes circulating in the microresonator. The CEO frequency is measured with an f-2f interferometer. Once the comb is stabilized, a close frequency comb tooth is beated with the 1550 nm source (The mode spacing of the frequency comb can be chosen so that there is a comb tooth close to 1550 or 1064 nm).

[Do have a look at this and tell me if anything is blatantly incorrect.]

[DYM:

  • You can remove the "computer" blocks except where we explicitly need them.
  • Make it clear what signals are present where: either make the diagram bigger and write down the fields, or add extra lines side by side, color coded to mean different things
  • Make your filter block look like this image (minus all the text); it's the symbol we generically use to mean "filter"
  • Remove the error signal block
  • Figure out what electrical signals are present where (you don't have to figure out exact calibration yet, just write down things like "Vsin(omega1 t) + sin(omega2 t)"
    • check out LaTeXiT if you use a MAC, this can generate nicely formatted equations in image format to stick into diagrams
  • Add a write up of the questions I asked you in the last couple meetings, as well as however much you know about the answer]
Attachment 2: 7.19.13.docx
  811   Thu Jul 18 16:57:58 2013 DmassNotesMeetings 

Buy phoenix (old notes have info)

Task list while we wait for new suspension to get machined

 

Nic: temperature noise coupling - can we estimate the level?

Need 2nd PDHv2

Need to enable PDHv2 boostage

Haixing BIO box for TTL signals to PDHv2 - there may exist one already

Card for Cymac

Acoustic noise hunting

The laser itself

CHeck out what E.G. has done @ 40m for lasers and boxes - get SURF to do another one for us?

Write down what AM/PM is from current modulation. What is RIN induced from current modulation

  810   Fri Jul 12 23:57:45 2013 nicolasLaserCavityAbsorbance of 532 nm on cavity mirror surface (pdf)

Quote:

 see attachment

 PDF for posterity

Attachment 1: ELog.pdf
ELog.pdf ELog.pdf
  809   Fri Jul 12 20:01:17 2013 Steve Maloney, Nicholas Smith-LefebvreLaserCavityAbsorbance of 532 nm on cavity mirror surface

 see attachment

Attachment 1: ELog.docx
  808   Thu Jul 11 16:46:19 2013 nicolas, steveDailyProgressLab WorkCalibrated beat shift due to green laser blast

The calibration will be applied soon, but the blast you see is about 3.5mW, and the beat calibration is 0.2539kHz/mV (and the beat signal you see is in Volts)

Attachment 1: Screen_Shot_2017-09-22_at_1.18.29_PM.png
Screen_Shot_2017-09-22_at_1.18.29_PM.png
  807   Mon Jul 8 13:02:17 2013 DmassLaserTransfer FunctionsLaser Transfer Function

DMASS NEEDS TO LOOK AT THESE PLOT TITLES AND FIX THEM (10/1/13)

 

Add G(f) = the driver transfer function as well + link the data on teh svn somewhere

 

 

We took some measurements with Rich a while ago in an effort to answer the question "where is all our phase going?"

We locked the laser to the cavity, an added in a drive signal (swept sine) at the PDH input, using an HP4395. We took the B/A transfer function, where B was taken from the PDH mixer before the low pass (using RF electronics), and A was taken at the input to the laser diode driver. 

This gives us the: "diode driver + laser + cavity + RFPD + mixer + time delay" transfer function (call this H(f))

We also took a transfer function straight through the diode driver using a 50 ohm terminator as a load (100 mA into 50 Ohms = 0.5 Watts), and putting the beefiest capacitor Rich could find across it as an AC coupling (call this G(f)). The pole from that (Z_in = 50 ohms + capacitor) was below the start frequency of the measurement.

I divided the diode driver transfer function out of the first transfer function (H(f)/G(f)). This yields:

"laser + cavity + RFPD + mixer + time delay",

so long as we assume that loading the diode driver with a 50 Ohm resistor didn't change its transfer function in any meaningful way (Rich assured us that because the diode driver was a Howland current source, it should not matter if we load it with a diode or a 50 Ohm terminator)

 

The first attached plot is the transfer function H/G, with a pole + time delay overplotted. The data points to there being another pole around 140 kHz (roughly the 3 dB point between the green and blue traces)

 

The 2nd plot is cavity pole + mystery pole @ 140 kHz + time delay. This happens to fall too fast. Can whatever looks like zeros up above 1MHz (if they are in fact zeros) pull the magnitude response up fast enough to make the green and blue overlap, or is this a clue that A) the transfer function is somehow not described well by simple poles/zeros or B) something is screwy with the measurement?

Attachment 1: laserTF1.png
laserTF1.png
Attachment 2: laserTF2.png
laserTF2.png
  806   Thu Jul 4 15:22:19 2013 nicolasLaserControl SystemRough tuning of laser temperature offset

We've noticed that sometimes the cavities prefer to have their PDH control signal to be offset slightly from zero mean, by where the temperature knob is set. I don't really know the reason for this, it doesn't seem like the actuator is hitting the rail. One known difference between the two cavities is that the west is still using the blue thor labs servo box.

The PDH control offload servo in the Cymac allows for an offset setpoint. I tuned the setpoints of the two cavities roughly to produce the lowest transmitted RIN of the cavities. I was able to reduce the fluctuations by an order of a few for the west cavity, while the easy cavity already seemed to be doing well at zero offset. The transmission of the west cavity also increased by about 8% with the new setpoint.

The attachment shows the cavity trans spectrum with and without the offsets. There is also a difference in the RIN of a factor of 3 between the cavities, this should be investigated. (Alignment?)

Here are the current settings:

controls@gaston:~$ ezcaread X1:CRY-W_LASER_SETPOINT
X1:CRY-W_LASER_SETPOINT = 400
controls@gaston:~$ ezcaread X1:CRY-E_LASER_SETPOINT
X1:CRY-E_LASER_SETPOINT = 0

Attachment 1: cavtrans.png
cavtrans.png
  805   Thu Jul 4 14:23:26 2013 nicolasLab InfrastructureGeneralCCD Camera power supply equipment and laser pointers in blue cabinet

We have a bunch of laser pointers of all colors, as well as many camera power supplies / power supply splitters in the blue cabinet.

I can't find the second camera that we plan to use for the REFL beams.

  804   Tue Jul 2 15:15:49 2013 DmassCryostatCavityNew Cavity Suspension

Quote:

 We want to redesign the cavity suspension because there is too much acoustic noise limiting the beat signal.

Nic and myself talked for a while about what the design should look like, and tried to compile a list of thoughts and questions. I made some sketches in my lab book, and showed them to Warren.

We have 7" of room vertically inside the experimental chamber.

The basic idea: use some sort of spring (BeCu coil springs, blade+wire) to suspend a secondary platform below the cold plate (~6" below), something like 6"x6" square. Rigidly mount the radiation shields to this platform, replacing the old flaky  thumb screw suspension. Make the setup modular enough so that if we want to iterate how we hang the cavity, we can do so easily. Use eddy current damping.

Questions we think we need to answer:

  • How soft / stiff of a spring do we want? Making this softer lowers the first resonance, and buys us more filtering. What are the other consequences of this?
  • What Q do we want?
  • What Q can we get?
  • How should we adjust pitch/yaw/height? (Nic suggested a weight that we could slide around)
  • How much contraction / throw do we get cooling from 300K to 120K by changing spring constant (Warren says ~10%. This is a large fraction of the clear aperture).
  • What do we want the thermal mass of the radiation shields to be? (Evan has been thinking nonstop about this, maybe he knows) - they are currently ~1/8" thick, maybe they can be thinner?
  • If we use coil springs, how should we attach them to the cold plate / platform?
  • What should the platform be made out of?

Warren also had an idea for how to make a coil spring out of a twisted pair of spring metals, one of which was coated in formvar, such that we got constrained layer damping

Warren:

make inertial damper! Use loudspeakers: Electromagnet for eddy current damping. Voice coil /  magnet geometries from loudspeakers. Add mass to voice coil, run resistor across voice coil. 

What mass do we need to make the inertial dampers?

Critical damping is what we want?

Rare earth magnets on bottom of exp chamber for ECD?

GS13 seismometers - look inside it to see how they suspend the spring? Talk to brian lantz see if he has pictures? Talk to Norna? Alastair? Callum?

Meeting time not good for Warren - reschedule regularly?

  803   Tue Jul 2 03:56:12 2013 DmassCryostatCavityNew Cavity Suspension

 We want to redesign the cavity suspension because there is too much acoustic noise limiting the beat signal.

Nic and myself talked for a while about what the design should look like, and tried to compile a list of thoughts and questions. I made some sketches in my lab book, and showed them to Warren.

We have 7" of room vertically inside the experimental chamber.

The basic idea: use some sort of spring (BeCu coil springs, blade+wire) to suspend a secondary platform below the cold plate (~6" below), something like 6"x6" square. Rigidly mount the radiation shields to this platform, replacing the old flaky  thumb screw suspension. Make the setup modular enough so that if we want to iterate how we hang the cavity, we can do so easily. Use eddy current damping.

Questions we think we need to answer:

  • How soft / stiff of a spring do we want? Making this softer lowers the first resonance, and buys us more filtering. What are the other consequences of this?
  • What Q do we want?
  • What Q can we get?
  • How should we adjust pitch/yaw/height? (Nic suggested a weight that we could slide around)
  • How much contraction / throw do we get cooling from 300K to 120K by changing spring constant (Warren says ~10%. This is a large fraction of the clear aperture).
  • What do we want the thermal mass of the radiation shields to be? (Evan has been thinking nonstop about this, maybe he knows) - they are currently ~1/8" thick, maybe they can be thinner?
  • If we use coil springs, how should we attach them to the cold plate / platform?
  • What should the platform be made out of?

Warren also had an idea for how to make a coil spring out of a twisted pair of spring metals, one of which was coated in formvar, such that we got constrained layer damping

  802   Tue Jul 2 03:39:50 2013 DmassNoise HuntingGeneralAcoustic Noise Hunting (part 1)

Quote:

 We borrowed the PZT noise hunting buzzer wand from the 40m (it is a bronze or copper dowel attached to a PZT with a BNC input).

Koji was kind enough to loan us a HV amp from the OMC lab (Thorlabs model MDT694A) - we are using it in the 100V max output mode (gain = 10, +/- 10V input, +/- 100V output).

I played around with it a bit yesterday with Koji and even with just the function generator (DS345), we could pretty clearly feel the buzz from 20 Vpp at 100-300 Hz.

Since we want to do band limited white noise, we ended up hooking up the DAC from the CyMac to the HV amp, and using awggui to make noise.

We put the DFD output into an audio receiver and listened to it on headphones while poking things with the calibrated poke stick.

At a signal of ~100V rms into the PZT, 

I turned down the noise level to the PZT wand until I could touch it to the table without hearing any audible difference in the beat noise on headphones. Excitation strength normalized between the different points by letting the weight of the PZT wand itself provide the force pushing the wand to the optic mount (or whatever I was touching).

The loudest couplings by were (in order of noisiness):

  1. Everything connecting to the diode (every part of the LM14S2, the cables connected to it, where the sidebands come in, and the butterfly package itself)
    1. This may be a consequence of the FM sidebands we are putting on, and getting acoustically driven perturbations to the 
  2. The hoses and flanges on the top of the cryostat which lead to the cold chamber
  3. The cryostat itself (poking the side) - this coupling was a good bit quieter than the flange on top of the cryostat which leads to the cold chamber

A couple steering mirrors (and the PDs) things in the PDH path and the readout path were audible as well, but these were both quieter than the butterfly mount and cryostat

Notes and thoughts:

  • The headphone noise sounded different (more high frequency content) when poking the readout optics compared to when we were poking the cryostat -> [ b/c phase vs frequency noise? ]
  • Coupling into cavity length couples directly into frequency
  • Readout is a phase detection

Another thing I noticed while listening to the beat  - when I was tuning the temperature of the laser via the knob on the front of the Thorlabs ITC510s in order to zero out the control signal to the laser, the noise would change a depending on what the control signal level was. The minimum noise (by ear) was not zero, and lock broke at different values on the + vs the - end (was asymmetric by somewhere between a factor of 1.5 and 2). Further investigation needed. I'm guessing it's either an error point offset, or us seeing some of the 7th order mode when we are slightly off one side of the cavity resonance

 Acoustic coupling: the top noise sources and our plans for mitigation:

Butterfly mount package

We are putting on sidebands at the diode itself via SMA connector on a PCB which touches the pins of the LD's butterfly package. We may be seeing sideband frequency modulation, or we may be seeing direct laser frequency noise. If the effect is from the sidebands, going to an EOM may fix this. Making little boxes for each diode is probably necessary: we could lift it off the table (via a rubber base?) and seismically isolate it. We also can thermally/acoustically isolate it, in case the coupling is from window / air currents / etc.

Cryostat flanges

The flanges are tired directly to the cold plate, which is tied directly to the suspension/frame/rad shield. Based on elog:787, I think that the stuff from ~200Hz -> 2kHz is the cavity support structure (it goes away when I let all the LN2 boil off). We are designing a v2 of the cavity shield support. The idea is to have ~four 6" springs hanging from the cold plate, suspending a 2nd plate, and mounting the cavities on top of that plate in something that looks very similar to their existing shields. There is a list of questions we need to answer in redesigning this.

Beat readout steering mirrors / beat PD / beamsplitter

The acoustic enclosure box we are making

PDH path optics

Ghost beams? Scattering? The first thing to try once we get to this is looking at the noise spectra and trying to dump the ghost beams to see if we can eliminate any scatter paths. This is in practice slightly difficult because it is so hard to see the ghost beams. Even with an IR card + IR camera, we expect not to be able to see most ghost beams.

 

  801   Sat Jun 29 13:09:54 2013 nicolasMiscControl Systemlaser temperature loops tuned

 

I tuned up the loops for the PDH control-offload to cavity temperature.

I also made a transmitted light lock threshold monitor that automatically shuts off the integrators (zero to filter bank input) when the cavity loses lock.

Attached is the time series of the control signal (going to laser temp) and error signal (control from PDH) for both cavities when the loops are closed.

 

Attachment 1: timeseries.pdf
timeseries.pdf
  800   Sat Jun 29 12:37:12 2013 nicolasComputingControl Systemsingle line command to rebuild x1cry model

I made an alias on the cymac to compile/install/restart the x1cry model in one command.

Just run "buildx1cry"

  799   Thu Jun 27 16:30:42 2013 DmassNoise HuntingGeneralAcoustic Noise Hunting (part 1)

 We borrowed the PZT noise hunting buzzer wand from the 40m (it is a bronze or copper dowel attached to a PZT with a BNC input).

Koji was kind enough to loan us a HV amp from the OMC lab (Thorlabs model MDT694A) - we are using it in the 100V max output mode (gain = 10, +/- 10V input, +/- 100V output).

I played around with it a bit yesterday with Koji and even with just the function generator (DS345), we could pretty clearly feel the buzz from 20 Vpp at 100-300 Hz.

Since we want to do band limited white noise, we ended up hooking up the DAC from the CyMac to the HV amp, and using awggui to make noise.

We put the DFD output into an audio receiver and listened to it on headphones while poking things with the calibrated poke stick.

At a signal of ~100V rms into the PZT, 

I turned down the noise level to the PZT wand until I could touch it to the table without hearing any audible difference in the beat noise on headphones. Excitation strength normalized between the different points by letting the weight of the PZT wand itself provide the force pushing the wand to the optic mount (or whatever I was touching).

The loudest couplings by were (in order of noisiness):

  1. Everything connecting to the diode (every part of the LM14S2, the cables connected to it, where the sidebands come in, and the butterfly package itself)
    1. This may be a consequence of the FM sidebands we are putting on, and getting acoustically driven perturbations to the 
  2. The hoses and flanges on the top of the cryostat which lead to the cold chamber
  3. The cryostat itself (poking the side) - this coupling was a good bit quieter than the flange on top of the cryostat which leads to the cold chamber

A couple steering mirrors (and the PDs) things in the PDH path and the readout path were audible as well, but these were both quieter than the butterfly mount and cryostat

Notes and thoughts:

  • The headphone noise sounded different (more high frequency content) when poking the readout optics compared to when we were poking the cryostat -> [ b/c phase vs frequency noise? ]
  • Coupling into cavity length couples directly into frequency
  • Readout is a phase detection

Another thing I noticed while listening to the beat  - when I was tuning the temperature of the laser via the knob on the front of the Thorlabs ITC510s in order to zero out the control signal to the laser, the noise would change a depending on what the control signal level was. The minimum noise (by ear) was not zero, and lock broke at different values on the + vs the - end (was asymmetric by somewhere between a factor of 1.5 and 2). Further investigation needed. I'm guessing it's either an error point offset, or us seeing some of the 7th order mode when we are slightly off one side of the cavity resonance

  798   Tue Jun 25 22:19:42 2013 ranaElectronicsLab WorkUSB Microscope: DigiMicro 2 MegaPixel

 Photo_on_6-25-13_at_10.17_PM.jpg

Works by connecting to Mac USB port and selecting it through PhotoBooth. No extra drivers needed.

  797   Tue Jun 25 15:30:14 2013 DmassNotesMeetingsMeeting Notes

Quote:

 Notes - Thur Jun 6 

Mike Martin's thesis for clock stability

Rana concerned about how high the broadband frequency noise is for the cavity
Suspended platform for cavities - one platform, both cavities on it
Coil springs for suspendoing the cavities?
Spring material - ask Warren (BeCu?)
f_res ~ 1/(2*pi) *sqrt(K/m_tot)
when we use rubber as a spring we get 1/f instead of 1/f^2 sooner (because Q is higher) - Q determines where the corner freq where 1/f^2 goes to 1/f 
Check the springs that Rana got from McMaster for Tara/PSL
Low frequency noise way too high?
New shield - black on the inside - can try black shields and shiny shields, and compare in situ
Chase the lower frequency noise
Parse Matt Arrans result into a transfer function - expect strong radiative
Grab PZT buzzer from 40m - skinny metal stick + pzt + cable to test mounts - use this for hunting which mounts are bad  (can one make a local sound wave gun?)
Start spring suspension ball rolling - check how much room inside experimental chamber - get extra shields machined
Consider the weight of the shields
Send shield assembly task to SURF
Talk to Huan! - have him help me understand past present and future of ground state cooling (and eventually frequency splitting)

 Tuesday 6/25/13

Described temperature laser modulation plans - Rana encouraged us to go to current modulation asap

PZT wand obtained - need voltage to drive it - PA85 boards in E shop (bud board) - hunt for HV amps - use band limited noise source

Buy blackout fabric / board from thorlabs

Start dumping beams

Evan is working on the temperature servos (sensing + actuation) 

Nic will buy free space AOMs from NEOS

 

  796   Mon Jun 24 17:23:59 2013 DmassDailyProgressstuff happensProgress

I made a cable to get the 2nd gold PD (which Rana so benevolently fixed for us) up and running;

Both cavities are now locked with the Gold PDs!

 

We made some progress on the temperature actuation front as well - We put a mirror by the input beam and bounced a 40 mW  405 nm laser off its edge to see if we could heat the cavity.

We also re-arranged the furniture when we were putting the lab back together, and moved the CyMac where the blue cryo parts cabinet used to be.

I got the beat realigned (we took it apart when trying to find a place to inject the heating beam), and the DFD signal back into the cymac

 

I turned the laser pointer on and off, there was a BOOMING SUPER FAST RESPONSE in the cavity cavity beat. The time constant at room temperature seems to be ~5 seconds (eyeballed in StripTool from the rate at which the beat signal asymptotes after giving it a heat pulse)

GOOD NEWS! We have a "way too much power" problem. This is much better than the opposite problem, and very promising for the prospect of using 

 

Next up:

We use the sweet sub Hz chopper that Steve brought in, and find out how small of a signal we can put in yet still extract via the spectrum (I guess we will want SNR(P_in), and for the final version SNR(P_in,T))

  795   Mon Jun 24 16:15:08 2013 nicolasComputingGeneralssh into gaston

I have the router forwarding port 2222 to port 22 on gaston.

So to ssh directly to gaston, do:

ssh controls@cymac1.caltech.edu -p 2222

  794   Mon Jun 24 15:46:42 2013 nicolasComputingDAQtaking CyMAC off the internet

I disconnected the ethernet cable going into the eth0 port of the cymac so that it's only connection is behind the router.

I had to change the EPICS environment variables on the cymac so that it didn't look for channels on the un-connected interface.

In files: /mnt/sshfs/cymac-tst/x1/target/fb/start_daqd.inittab (for the daqd process) and
/opt/rtcds/rtcds-user-env.sh for user logins.

I edited the /etc/network/interfaces file on the cymac to no longer set the eth0 interface to static.

Now the router at 10.0.5.1 assumes the static ip for cymac1.caltech.edu and forwards ports 22 and 8088 to the cymac.

  793   Fri Jun 21 11:51:06 2013 ranaComputingDAQWorkstation setup wiki

  Info from K Thorne on how to set up our CDS workstations:

https://awiki.ligo-wa.caltech.edu/aLIGO/CdsUbuntu12ToolBuild

  791   Thu Jun 20 15:41:46 2013 Steve, nicolas, d-massDailyProgressLab WorkAttempt at cavity heating with laser pointer

While looking at the PDH control signal (our length sensor) we shot green and red laser pointers into the cavity through the diaphram. We shot the lasers through the transmission side using the west cavity.  5 mWatt lasers failed to give a distinctive, readily identifiable length change that gave us any degree of confidence on the cause of the change. At his time we cannot attribute length changes to energy input from laser.  One possible solution may be to use a more powerful laser. 

Dmass also suggests using a beam chopper to get some time modulation.

  790   Thu Jun 20 11:32:59 2013 DmassNotesMeetingsMeeting Notes

 Notes - Thur Jun 6 

Mike Martin's thesis for clock stability

Rana concerned about how high the broadband frequency noise is for the cavity
Suspended platform for cavities - one platform, both cavities on it
Coil springs for suspendoing the cavities?
Spring material - ask Warren (BeCu?)
f_res ~ 1/(2*pi) *sqrt(K/m_tot)
when we use rubber as a spring we get 1/f instead of 1/f^2 sooner (because Q is higher) - Q determines where the corner freq where 1/f^2 goes to 1/f 
Check the springs that Rana got from McMaster for Tara/PSL
Low frequency noise way too high?
New shield - black on the inside - can try black shields and shiny shields, and compare in situ
Chase the lower frequency noise
Parse Matt Arrans result into a transfer function - expect strong radiative
Grab PZT buzzer from 40m - skinny metal stick + pzt + cable to test mounts - use this for hunting which mounts are bad  (can one make a local sound wave gun?)
Start spring suspension ball rolling - check how much room inside experimental chamber - get extra shields machined
Consider the weight of the shields
Send shield assembly task to SURF
Talk to Huan! - have him help me understand past present and future of ground state cooling (and eventually frequency splitting)
  789   Wed Jun 19 11:41:11 2013 DmassNoise HuntingNoise BudgetBeat Readout Box

The 40m has this incredibly nice Cadillac of a box on the end table which Steve designed.

  1. I talked to Steve, looked at the box, had some ideas on how to keep it simple and cheap:
    1. No handles
    2. No crazy coating/films
    3. Simpler version of o-ring which doesn't require machining race-track groove on top and bottom edges of box
    4. Copy the 40m "no windows" design. In our case between the box and the cryostat (connect to the cryostat instead - these can be added in a later generation if we want them) - connecting the box to the cryostat seems much simpler and cheaper than making windows
  2. Found out what plastic / connectors they used
    • 1" thick acrylic plastic (McMaster #8560K328)
    • Expansion inserts specifically for plastics (McMaster #93415A071)
  3. I looked at the hole diameter required for the plastic inserts for the various screws, and what thickness of plastic that would require
    • 1/4-20 screws need 5/16 for insert, so 1" or 15/16" acrylic are the options that give us more than 3/16" clearance on a side (which is fairly thin - I wouldn't want to tear out a hole if I accidentally compressed the box)
    • #10 screws need 1/4" for insert, so 1" or 15/16" or 11/16" would probably work here
  4. I mocked up a drawing of what the box would look like (below), and checked to see whether we could get all the cuts we would need out of a 48"x48" sheet (the answer was yes)
  5. I checked how much the box would weigh if I made it out of 1" plastic:
    • Acrylic density: 0.48 lbs/in^3
      • http://www.engineersedge.com/wwwboard/posts/15096.html
      • http://www.usplastic.com/knowledgebase/article.aspx?contentkey=884
    1. Lid: 20 lbs
    2. Main box (w/o lid): ~40 lbs
  6. I decided to try to make it lighter to make it easier to manhandle (I wanted to be able to pick up the whole thing without its own weight putting a lot of stress on the edge connections / screws)
  7. Making the plastic thinner begins to exclude the use of screw sizes that seem reasonable (e.g. with 1/2" plastic it might be sketchy to put  a 1/4" insert and leave only 1/8" of material on each side if we start twisting the box)
  8. I searched for L-brackets instead, it feels like a sturdier solution (for the same reason you use L-brackets when framing a house instead of just end-nailing everything)
  9. I chose some L-brackets (and bought them):
    1. Style 9 and 13 from http://www.mcmaster.com/#corner-brackets/=n9ic1q ( 17715A44  and 15275A64, which have clearances for  #8 and 1/4" screws, respectively)
    2. Style 9 is prettier with smaller (and more) screw holes, I will use it if I can get away with #8 screws (probably not). Style 13 uses 1/4-20's, though I like the layout of the clearance a lot less
  10.  Finalized initial design so that I was satisfied that it would work
  11. Ordered:
    1. Flat foam ribbon for "gasket"/seal
    2. 48x48" sheet of 1/2" acrylic
    3. 1/4-20 L-brackets

I'm currently making drawings for the cuts they need to do in the shop, and will see if the guys that steve used have time.

I'm operating on a "minimize the amount of machining that needs to be done to make a functional acoustic enclosure" mission statement.

In other words: I designed a box based on the 40m box, changed some things, and bought the materials necessary for it to be built

Attachment 1: BeatReadoutBox.png
BeatReadoutBox.png
  788   Mon Jun 17 13:22:11 2013 nicolasMiscControl SystemTemperature response of west cavity

Here is a transfer function of the temperature control plant of the west cavity.

A 2Hz loop should be possible with the amount of delay seen here.

data in ~/CryoLabSVN/Measurements/tempFeedback/2013-06-17

Attachment 1: westcav_TF.pdf
westcav_TF.pdf westcav_TF.pdf
  787   Thu Jun 13 01:39:23 2013 DmassUpdateLab WorkComparison of beats

 [6/18/13 - dym expanded this syncopation of an elog]

One of the things we want to know, and are fairly uncertain about, is: "How much temperature and acoustic noise does the LN2 boiling create, and how much does this couple into our measurement?"

In an attempt to answer this question, I let the LN2 boil off completely, and then made some beat note measurements. We have a fairly sizable window of time after all the LN2 boils off where the cavity starts to *slowly* warm up. The procedure for that:

  1. Let LN2 boil off
  2. Valve off vacuum chamber
  3. Turn off turbo pump (it creates far too much vibrational noise to make low noise beat measurements)
  4. Measure beat note
  5. Monitor pressure, periodically pump down if the pressure starts climbing too high (either via leaks or heating up molecules which were previously cryopumped onto some surface inside the cryostat)
  6. Repeat 2-4

 While we were doing this, we also made progress getting interesting signals into the CyMac.

===============================================================

Cymac progress:

  1. We plugged the following signals into the CyMac:
    • Beat signal! (DFD output) 
    • East cavity transmission
    • West cavity transmission
    • East laser PDH control signal
    • West laser PDH control signal
  2. And take the following signals out of the CyMac:
    • Temp feedback to East cavity temperature controller (ITC510)
    • Temp feedback to West cavity temperature controller (ITC510)
  3. Because there was no anti aliasing, I made a little board with BNC inputs and outputs with an 8 kHz low pass in each channel
    • 200 Ohm resistor, 100 nF capacitor
  4. I put the beat, the transmission, and the control signals through this.
  5. I have done nothing to mitigate anti imaging, and don't understand it except from a "timer reversal symmetry of anti aliasing" perspective
  6. We looked at the DC level going into the CyMac on the beat channel to get the calibration (recorded as a gain in the model - currently offline due to cleaning)
  7. We tested the noise of each channel by putting a terminator on the outside of the makeshift AA filter and looking at the noise spectrum
    • One of the channels had really high noise (?check? maybe West cavity transmission) - it went down when I wiggled the relevant clippy doodad
    • The noise of the channels were close to what I though they should be - in the old test stand with the old AA chassis, the noise was ~6 uV/rtHz. The noise measured using the non-ideal setup was ~3 uV/rtHz. 

===============================================================

Attachment 1: BeatComparisons_A.png
BeatComparisons_A.png
  786   Mon Jun 10 17:46:21 2013 nicolasUpdateLab WorkToday's measurements in a local SVN checkout

I checked out the cryolab svn directory to ~/CryoLabSVN .

Todays measurements are in ~/CryoLabSVN/Measurements/CavityBeat125/2013-06-10 .

  785   Sun Jun 9 03:32:37 2013 ranaElectronicsGeneralRFPD ready - s/n #116

 Finally found the issue with this RFPD: there was in internal short from one of the legs of the main tank circuit inductor to ground. After removing this leg from the board and jumpering it over with a flying trace, all seems to function as it should.

Pictures of the new circuit as well as screenshots of the analyzer during the measurements and the marked up schematic are on the picasa site here.

RFPD 116 is on my desk in the 40m and ready to be taken and used. Analysis of data to be posted here next.

  784   Fri Jun 7 12:55:00 2013 nicolasLab InfrastructureRoadmapautomation processes

As a summer project, a high school student might work on the following automation tasks:

  • A monitor of cavity temperature and pressure that sends us an email when they go outside of nominal values.
  • Frequency counter GPIB script that puts the absolute beat frequency into EPICS.
  • more?
  783   Thu Jun 6 16:53:50 2013 nicolasElectronicsDAQADC channels for signals

When Dmass is going to plug signals into the ADC he should do this:

BEAT -> ADC24
East Trans -> ADC25
West Trans -> ADC26

The channel names will be:

X1:CRY-BEAT_MON_*
X1:CRY-E_TRANS_*
X1:CRY-W_TRANS_*

These are not represented on the medm screens yet.

  782   Thu Jun 6 16:44:21 2013 nicolas, dmassDailyProgressControl Systemslow DIGITAL laser temp control

Today we set up slow feedback of the PDH control signal to the laser temperature for both cavities. The control filtering all happens in the cymac.

We have some new screens to help with this.

Both loops are closed with an integrator and about 1Hz BW.

laser_temp_feedback.png

Attachment 1: laser_temp_feedback
  781   Thu Jun 6 15:03:41 2013 nicolasComputingDAQThe state of gaston

DTT now can get testpoint data, and the synchronization errors have gone away.

 I am not sure why the sync errors went away, i did install ntp while trying to fix something else so maybe that updated our clock to be close enough to the cymac.

The testpoint problem was fixed with the help of Chris Wipf. There was some trouble with the DTT RPC communication with the awgtpman processes running on the cymac. The command "diag -i" tells us something about what DTT thinks about the network configuration. It was giving this before:

controls@gaston:~$ diag -i
Diagnostics configuration:
awg 5 0 127.0.1.1 822095877 1 10.0.5.11
awg 7 0 127.0.1.1 822095879 1 10.0.5.11
awg 8 0 127.0.1.1 822095880 1 10.0.5.11
nds * * 10.0.5.11 8088 * 127.0.0.1
tp 5 0 127.0.1.1 822091781 1 10.0.5.11
tp 7 0 127.0.1.1 822091783 1 10.0.5.11
tp 8 0 127.0.1.1 822091784 1 10.0.5.11

The problem turned out to be all those 127.0.1.1 lines. My guess is that the problem is new because we have our framebuilder and nds running on the same machine as our frontend. So when the nds tries to get the ip address of the frontend (where awgtpman is running), the host lookup for cymac1 returns 127.0.1.1, instead of the correct ip address.

The fix is a bit of a hack, but I edited the hosts file (of the cymac) so that cymac1 points to 10.0.5.11, instead of 127.0.1.1. Now diag -i gives:

controls@gaston:~$ diag -i
Diagnostics configuration:
awg 5 0 10.0.5.11 822095877 1 10.0.5.11
awg 7 0 10.0.5.11 822095879 1 10.0.5.11
awg 8 0 10.0.5.11 822095880 1 10.0.5.11
nds * * 10.0.5.11 8088 * 127.0.0.1
tp 5 0 10.0.5.11 822091781 1 10.0.5.11
tp 7 0 10.0.5.11 822091783 1 10.0.5.11
tp 8 0 10.0.5.11 822091784 1 10.0.5.11

With this configuration, DTT can get testpoints.

I am not sure if there is a better solution than messing with the hosts file. It is not great because it will have to change if the network is reconfigured.

  780   Wed Jun 5 02:17:30 2013 DmassNoise HuntingNoise BudgetHunting for intensity noise coupling

Quote:

 ** do I need to divide by the BW of my measurement to get my answer into PSD/ASD?

 YES, YES I DID!

I played with the 785 and convinced myself I understood what it was doing w.r.t. bin width, and getting from dBVpk-> Vrms/rtHz.

I calibrated all the plots I made and attached them here as one pdf.

There are two more transfer function plots which don't display properly on the thumbnail, but if you click the image they load fine.

I had to pay attention to calibration because the beat calibration also changes as I change the transmitted power levels.

Maybe I should change this into "Hz/(Intracavity) Watts" to normalize it more intelligently?

Attachment 1: AllPlots.pdf
AllPlots.pdf AllPlots.pdf AllPlots.pdf AllPlots.pdf AllPlots.pdf AllPlots.pdf AllPlots.pdf AllPlots.pdf
  779   Tue Jun 4 11:38:59 2013 nicolasComputingDAQThe state of gaston

Gaston currently has launcher icons for:

Dataviewer (and it works now)
MEDM
StripTool
and DTT

There is a problem with DTT right now, usually gives a synchronization error (though if you do a long time series measurement of a DQ channel it works). It also gives an error on launch saying that testpoints and awg are not supported. Still needs some debugging.

  778   Tue Jun 4 03:30:53 2013 DmassNoise HuntingNoise BudgetHunting for intensity noise coupling

An attempt at documenting and/or explaining where I am stuck/confused in my intensity noise coupling measurement.

After a long night of measuring, I realized I had taken the data in dBVpk .

 

 

I also took some transfer functions, but until I trust my calibrations I am going to forego posting dB's which mean nothing.

THIS is a spreadsheet which contains info about the various measurements, and applicable calibrations.

The last plot (calibration) is straightforward to interpret. I spun the wheel on the marconi while driving the DFD, and wrote down DC voltages coming out. It has a slope of 1.14e6 Hz/V

To convert from dBVpk to Vrms I did:

Vrms = 10^(dBVpk/20)/sqrt(2)

I then applied the calibration of 1.14e6 Hz/V, and got a spectrum that was a factor of 20 too high.

I went through my lab book and my code for both this and the spectrum I posted last week, but cannot find errors, just inconsistencies.

 

** do I need to divide by the BW of my measurement to get my answer into PSD/ASD?

Attachment 1: EastMeasAlign.pdf
EastMeasAlign.pdf
Attachment 2: EastMeasMisalign.pdf
EastMeasMisalign.pdf
Attachment 3: WestMeasAlign.pdf
WestMeasAlign.pdf
Attachment 4: WestMeasMisalign.pdf
WestMeasMisalign.pdf
Attachment 5: Calibration.pdf
Calibration.pdf
  777   Tue Jun 4 03:22:56 2013 DmassNoise HuntingNoise BudgetHunting for intensity noise coupling

I spent some of last week trying to get a measurement for the intensity noise coupling.

I have PDs in transmission on:

  1. The Beat
  2. East Cavity Intensity
  3. West Cavity Intesity 

I measured the following quantities for different misalignments of the input beam's pitch

  • Coherence between intensity and beat for each cavity
  • Beat Freq Noise Spectra
  • Intensity Noise Spectra
  • DC Transmitted power for east/west cavities

I used the DC transmitted power to parametrize the input misalignment

 

I am currently plotting things and doing sanity checks: "<Make sure the level of the calibrated beat signal is consistent with previous measurements"

I am failing this sanity check, and this beat is appearing a factor of 10-20 higher than previous measurements. Trying to figure out what the heck is going on.

  776   Tue Jun 4 03:22:55 2013 DmassNoise HuntingNoise BudgetHunting for intensity noise coupling

I spent some of last week trying to get a measurement for the intensity noise coupling.

I have PDs in transmission on:

  1. The Beat
  2. East Cavity Intensity
  3. West Cavity Intesity 

I measured the following quantities for different misalignments of the input beam's pitch

  • Coherence between intensity and beat for each cavity
  • Beat Freq Noise Spectra
  • Intensity Noise Spectra
  • DC Transmitted power for east/west cavities

I used the DC transmitted power to parametrize the input misalignment

 

I am currently plotting things and doing sanity checks: "<Make sure the level of the calibrated beat signal is consistent with previous measurements"

I am failing this sanity check, and this beat is appearing a factor of 10-20 higher than previous measurements. Trying to figure out what the heck is going on.

  775   Mon Jun 3 14:50:24 2013 nicolasComputingControl Systemmedm screens

I set up autofs/sshfs on gaston to mount the/opt/rtcds/tst directory over the network according to this guide.

This way the autogenerated screens are accessible on the workstation.

_____

So now the accepted way to work on screens is to:

edit your user created screens in /opt/rtcds/userapps/CyMAC/medm_screens (on gaston)

svn add any new screens

Run $sync_cymac "your commit message"

This will do a svn commit, and a svn update on the cymac.

I think links to autogenerated screens should be absolute path, this is OK as long as the path to the autogenerated screens is mounted as I did above. If we want to start running medm on remote laptops, we may want to think about doing relative paths.

____

Then you can see your screens by clicking the MEDM launcher on the desktop, or running the command 'sitemap' in a terminal.

  774   Sun Jun 2 11:03:41 2013 illustratorComputingDAQhow i got dataviewer running on gaston (the workstation)

Walt-Disney-Screencaps-Gaston-walt-disney-characters-31944510-2560-1440.jpg?

  773   Sat Jun 1 17:47:39 2013 ranaComputingDAQNDS1 Proxy

 

 None of our Cymacs should be directly accessible from the outside - not safe.

  772   Thu May 30 19:36:33 2013 nicolasComputingDAQNDS1 Proxy

Quote:

The NDS1 Proxy software is in

/cvs/cds/caltech/target/ndsproxy

on nodus. I think that if its run on the Cymac/FB it should open up a port by which we can use DTT, DV, and get_data, and LIGO_DC to get data, etc.

 This is probably only necessary if you can't see the NDS server directly, like if it's behind a firewall.

In our case, the workstation is on the same network as the cymac, and if we really wanted data over the internet, the cymac itself is also facing directly to the internet.

  771   Thu May 30 19:35:10 2013 nicolasComputingDAQhow i got dataviewer running on gaston (the workstation)

Dataviewer now runs directly on gaston without having to crap around with sshing into the cymac.

open a terminal and run "dataviewer -s cymac1"

It doesn't have the graphical and segfault bugs that the cymac version has.

____

First I untared dv and grace from here. It all lives in /ligo/apps/ubuntu12/

I had to install the libmotif4 package because of a missing .so

made symbolic links to all the grace and dv binaries in /ligo/apps/ubuntu12/bin and added this dir to the PATH variable in ~/.profile.

I also added a line for cymac1 in the /etc/hosts file.

I don't yet know how to make cymac1 the default NDS server when you open dataviewer. Right now it defaults to "-a". Setting the NDSSERVER environment variable didn't help.

Next I will try to get EPICS/MEDM/Striptool and DTT running.

 

EDIT: I also had to install the package xfonts-75dpi so that grace would stop giving an error

Also, my cymac hosts entry was backwards, so it was trying to connect to the cymac over the internet instead of the direct connection in the internal network. That's fixed now. I will make an icon in the ubuntu launcher for dataviewer.

ELOG V3.1.3-